OpenStudy (anonymous):

A 250 mL sample of nitrogen(N2) has a pressure of 745 mm Hg at 30 C.What is the mass of nitrogen?

OpenStudy (sriram):

use P*V=n*R*T we know P,V,R & T find n i.e. number of moles mass=n*14

OpenStudy (preetha):

Convert the mm Hg to atm first. 760 mm Hg = 1 atm

OpenStudy (preetha):

And C to Kelvin. K = C+ 273.15